OpenStudy (anonymous):

Through (1,6) and (5,2) use function notation

OpenStudy (alexandervonhumboldt2):

yeah so

OpenStudy (alexandervonhumboldt2):

find slope using slope formula (y1-y2)/(x1-x2)=(6-2)/(1-5)=4/-4=-1 so slope is -1

OpenStudy (alexandervonhumboldt2):

now substitue point in y=-x+b 2=-5+b b=????????

OpenStudy (anonymous):

sorry I stepped away for a moment.

OpenStudy (anonymous):

B=7

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

so the final equation is y=-x+7